# BaCo2(AsO4)2: Strong Kitaev, After All

### v3

@@ Abstract
Earlier such attempts that relied on the easy-plane ${\sf J}_{1}$${\sf J}_{2}$${\sf J}_{3}$ $XXZ$ model led to a conundrum: the extracted parameters that fit the high-field INS data did not correspond to the expected spiral ground state, nor was the model able to explain the observed fully gapped zero-field spectrum even when assuming such a state [9, 10].
A recent study provided more INS data [40], but rejected substantial Kitaev terms in BaCAO in favor of largely the same easy-plane ${\sf J}_{1}$${\sf J}_{3}$ $XXZ$ model, imposing the zero-field order as a classical spiral state [11]. This model yielded only a rough fit to a subset of the high-field INS data and could not reconcile puzzling low critical fields. The growing enigma became even more bewildering as recent theoretical works suggested a complete absence of the spiral phase in the phase diagram of the quantum $S\!=\!1/2$ ${\sf J}_{1}$${\sf J}_{3}$ $XXZ$ model [54, 55].
@@
Figure 1: Magnon energy vs ${\bf q}$ for ${\bf B}\!\parallel\!b$ and ${\bf B}\!\parallel\!a^{*}$; $B\!=\!3$ T. Diamonds [9], squares [40], and circles (this work) are the INS data, and solid lines are theory fits. $\Delta E_{\rm M}$ is the energy splitting at M1. Insets: Crystallographic axes and principal directions, 1st-, 2nd-, and 3rd-neighbor bonds on the honeycomb lattice with bond-dependent exchanges represented by colors, and the Brillouin zone with $\Gamma$ and M1 points indicated.
@@
This is the case with the prior INS high-field results for BaCAO [9, 10, 40]. All the existing data are for fields along one principal in-plane direction, ${\bf B}\!\parallel\!b$, perpendicular to the bond of the honeycomb lattice. Much of the results are for momentum along the high-symmetry $\Gamma$M1 line in the Brillouin zone (BZ), in the plane normal to the field; see insets in Fig. 1. This is precisely the direction for which the spectrum is independent of one of the key bond-dependent Kitaev-like terms.
Although the BaCAO spectra for other momentum directions for that field orientation were also probed, their analyses were either unaware of Kitaev terms [10] or did not consider these data separately from those in the $\Gamma$M1 direction to infer such terms [40]. In these studies, easy-plane $XXZ$ models were used
@@
where $\langle ij\rangle_{n}$ for $n\!=\!1,2,3$ are the $n$th-neighbor bonds, ${\sf J}_{n}$ are the exchanges (${\sf J}_{1,2}\!<\!0$, ${\sf J}_{3}\!>\!0$), and $\Delta_{n}$ are the $XXZ$ anisotropy parameters for such bonds. In BaCAO, ${\sf J}_{2}$ is weaker than ${\sf J}_{3}$, as is typical in the honeycomb-lattice magnets [10, 17], and was neglected in Ref. [40]. All $\Delta_{n}$ are easy-plane-like, ${\bf H}\!=\!\mu_{B}{\bf B}$, and the g-tensor is diagonal, $\mathbf{\hat{g}}\!=\!\{g_{x},g_{y},g_{z}\}$, with $g_{x}\!\approx\!g_{y}\!\approx\!5.0$ and $g_{z}\!\approx\!2.5$ [10], where we use crystallographic frame in which $x\!=\!a^{*}$ and $y\!=\!b$; see inset of Fig. 1.
@@
Our Figure 1 presents key and unambiguous evidence that the description of BaCAO with model (1) is incomplete and that very significant bond-dependent terms must be present in it. Diamonds and squares in Fig. 1 are the prior data from Refs. [9] and [40], respectively, for the spin-flip energy along the $\Gamma$M1 line for the ${\bf B}\!\parallel\!b$ orientation and $B\!=\!3$ T, well above the critical field of $B_{c}\!\approx\!0.55$ T [10]. Crucially, Fig. 1 presents previously unpublished results (circles), obtained at IN22 spectrometer at the Institut Laue Langevin, Grenoble, for the same momentum direction and field value, but for a different principal in-plane field orientation, ${\bf B}\!\parallel\!a^{*}$; see End Matter (EM) for the experimental details. Contrary to the expectations from the $XXZ$ model, the two bands for different field orientations show a large splitting.
@@
Figure 2: (a) The field dependence of $E_{\Gamma}$ (circles) and $E_{\rm M_{1}}$ (diamonds) for ${\bf B}\!\parallel\!b$ and ${\bf B}\!\parallel\!a^{*}$, shown as black and blue symbols (data) and lines (theory), respectively. $\Delta E_{\rm M}$ is indicated and shown in the inset. (b) and (c), Same as in Fig. 1: data and their fits for magnon energies vs ${\bf q}$ with ${\bf B}\!\parallel\!b$, $B\!=\!0.7$ T, 3.0 T, and 6.0 T from Refs. [9, 10] in (b), and with ${\bf B}\!\parallel\!a^{*}$, $B\!=\!0.6$ T, 1.0 T, 2.0 T, and 3.0 T (this work) in (c). Dashed lines represent the lowest edges of the two-magnon continua for 0.7 T in (b), 0.6 T in (c), and for the ${\rm M_{1}}$ point in (a).
@@
Figure 3: (a) Data from Ref. [40] for ${\bf B}\!\parallel\!b$, $B\!=\!3$ T, lines are as described in the legend. (b) Intensity maps of ${\cal S}({\bf q},\omega)$ from Ref. [40] and our results. Dashed lines mark the bottom of the two-magnon continuum. (c) The ${\sf J}_{z\pm}$${\sf J}_{3}$ phase diagram of the model (1)$+$(2) with the LT and DMRG boundaries and phases; $\Delta_{n}$, ${\sf J}_{2}/{\sf J}_{1}$, and ${\sf J}_{\pm\pm}/{\sf J}_{1}$ are from (3), symbols show various parameter sets (see the text). (d) Results from DMRG $12\!\times\!12$ cluster calculations for the set (3) with spins in the $yz$ plane.
In Fig. 2(b), we reproduce the INS data from Refs. [9, 10] for the magnon energies along the $\Gamma$M1 line for ${\bf B}\!\parallel\!b$ with $B\!=\!0.7$ T, 3.0 T, and 6.0 T. Fig. 2(c) presents previously unpublished results for ${\bf B}\!\parallel\!a^{*}$ with $B\!=\!0.6$ T, 1.0 T, 2.0 T, and 3.0 T, showing the field evolution of the band that remains much flatter than the one in Fig. 2(b). The solid lines represent the theoretical fits.
@@
It is the ${\sf J}_{z\pm}$ term in (2) to which the $\Gamma$M1 INS data for the high ${\bf B}\!\parallel\!b$ fields are insensitive, making it difficult to detect [53]. It is also the only term in the model (1)$+$(2) that can induce a tilt of spins out of the crystallographic plane, as it is the only one that couples the in-plane and out-of-plane spin components [18, 19].
@@
This is the set of parameters that provides a close fit to all available INS data for BaCAO in the polarized phase, as shown in Figs. 1 and 2, and Figs. 3(a) and 3(b) for ${\bf B}\!\parallel\!b$ at $B\!=\!3$ T with the data from Ref. [40]. In Fig. 3(a), we show the data in multiple ${\bf q}$-panels together with the $XXZ$ fit from Ref. [40] (black dashed lines) and our results for the single-magnon branches and the bottom of the two-magnon continuum (solid and light-blue dashed lines, respectively). Fig. 3(b) shows a side-by-side comparison of the intensity maps of the dynamical structure factor ${\cal S}({\bf q},\omega)$ from Ref. [40] and our theoretical results. The comparison of our results with the data for $B\!=\!0.75$ T and ${\bf B}\!\parallel\!b$ from Ref. [40] is shown in EM.
@@
In Fig. 3(c), we present the classical Luttinger-Tisza (LT) ${\sf J}_{z\pm}$${\sf J}_{3}$ phase diagram of the model (1)$+$(2) for $\Delta_{1}\!=\!0.36$ and ${\sf J}_{2}/{\sf J}_{1}\!=\!0.03$ as in (3), and we set $\Delta_{3}\!=\!{\sf J}_{\pm\pm}\!=\!0$ for simplicity [19, 64]. The sequence of the FM, ZZ, and helical phases shown in Fig. 3(c) is common in many classical honeycomb-lattice models [11, 54, 19].
@@
Recent studies of the quantum $S\!=\!1/2$ edition of the $XXZ$ ${\sf J}_{1}$${\sf J}_{3}$ model using the density-matrix renormalization group (DMRG) method [54, 55] found no trace of the spiral phase in its phase diagram, uncovering instead only narrow ranges of non-classical states, such as dZZ. Here, we extend this study to the ${\sf J}_{1}$${\sf J}_{z\pm}$${\sf J}_{3}$ model, with the remaining parameters ($\Delta_{n}$, ${\sf J}_{\pm\pm}$, and ${\sf J}_{2}$) fixed according to the proposed BaCAO set in (3).
@@
DMRG simulation of the proposed parameter set (3) on a $12\times 12$ cylinder are shown in Fig. 3(d), demonstrating the dZZ state with spins shown in the $yz$ plane to emphasize their out-of-plane tilt. While the non-classical dZZ state for the $XXZ$ ${\sf J}_{1}$${\sf J}_{3}$ models has recently received theoretical endorsements [42, 66, 58, 55], the tilt is unusual and is clearly due to the ${\sf J}_{z\pm}$ coupling. According to our DMRG results, every second spin in the dZZ structure is tilted out of plane by about $35\degree$. Using the g-factors $g_{z}/g_{ab}\!\approx\!0.5$, this yields $7.8\degree$ average tilt of the magnetic moment, compared with $\sim\!6\degree$ suggested by neutron polarimetry [12, 13]. These and other quantitative discrepancies can be attributed either to finite-size effects in DMRG or to small terms in the description of BaCAO that are not accounted for by our model [62].

### v5

@@ Abstract
Earlier such attempts that relied on the easy-plane ${\sf J}_{1}$${\sf J}_{2}$${\sf J}_{3}$ XXZXXZ model led to a conundrum: the extracted parameters that fit the high-field INS data did not correspond to the expected spiral ground state, nor was the model able to explain the observed fully gapped zero-field spectrum even when assuming such a state [9, 10].
A recent study provided more INS data [40], but rejected substantial Kitaev terms in BaCAO in favor of largely the same easy-plane ${\sf J}_{1}$${\sf J}_{3}$ XXZXXZ model, imposing the zero-field order as a classical spiral state [11]. This model yielded only a rough fit to a subset of the high-field INS data and could not reconcile puzzling low critical fields. The growing enigma became even more bewildering as recent theoretical works suggested a complete absence of the spiral phase in the phase diagram of the quantum S=1/2S\!=\!1/2 ${\sf J}_{1}$${\sf J}_{3}$ XXZXXZ model [54, 55].
@@
Figure 1: Magnon energy vs $\bf q$} for ${\bf B}\!\parallel\!b$ and ${\bf B}\!\parallel\!a^{*}$; B=3B\!=\!3 T. Diamonds [9], squares [40], and circles (this work) are the INS data, and solid lines are theory fits. $\Delta E_{\rm M}$ is the energy splitting at M1. Insets: Crystallographic axes and principal directions, 1st-, 2nd-, and 3rd-neighbor bonds on the honeycomb lattice with bond-dependent exchanges represented by colors, and the Brillouin zone with $\Gamma$ and M1 points indicated.
@@
[…]
@@
where $\langle ij\rangle_{n}$ for n=1,2,3n\!=\!1,2,3 are the $n$th-neighbor bonds, ${\sf J}_{n}$ are the exchanges (${\sf J}_{1,2}\!<\!0$, ${\sf J}_{3}\!>\!0$), and $\Delta_{n}$ are the $XXZ$ anisotropy parameters for such bonds. In BaCAO, ${\sf J}_{2}$ is weaker than ${\sf J}_{3}$, as is typical in the honeycomb-lattice magnets [10, 17], and was neglected in Ref. [40]. All $\Delta_{n}$ are easy-plane-like, ${\bf H}\!=\!\mu_{B}{\bf B}$, and the g-tensor is diagonal, $\mathbf{\hat{g}}\!=\!\{g_{x},g_{y},g_{z}\}$, with $g_{x}\!\approx\!g_{y}\!\approx\!5.0$ and $g_{z}\!\approx\!2.5$ [10], where we use crystallographic frame in which $x\!=\!a^{*}$ and y=by\!=\!b; see inset of Fig. 1.
@@
[…]
@@
Figure 2: (a) The field dependence of $E_{\Gamma}$ (circles) and $E_{\rm M_{1}}$ (diamonds) for ${\bf B}\!\parallel\!b$ and ${\bf B}\!\parallel\!a^{*}$, shown as black and blue symbols (data) and lines (theory), respectively. $\Delta E_{\rm M}$ is indicated and shown in the inset. (b) and (c), Same as in Fig. 1: data and their fits for magnon energies vs $\bf q$} with ${\bf B}\!\parallel\!b$, B=0.7B\!=\!0.7 T, 3.0 T, and 6.0 T from Refs. [9, 10] in (b), and with ${\bf B}\!\parallel\!a^{*}$, B=0.6B\!=\!0.6 T, 1.0 T, 2.0 T, and 3.0 T (this work) in (c). Dashed lines represent the lowest edges of the two-magnon continua for 0.7 T in (b), 0.6 T in (c), and for the $\rm M_{1}$} point in (a).
@@
Figure 3: (a) Data from Ref. [40] for ${\bf B}\!\parallel\!b$, B=3B\!=\!3 T, lines are as described in the legend. (b) Intensity maps of ${\cal S}({\bf q},\omega)$ from Ref. [40] and our results. Dashed lines mark the bottom of the two-magnon continuum. (c) The ${\sf J}_{z\pm}$${\sf J}_{3}$ phase diagram of the model (1)++(2) with the LT and DMRG boundaries and phases; $\Delta_{n}$, ${\sf J}_{2}/{\sf J}_{1}$, and ${\sf J}_{\pm\pm}/{\sf J}_{1}$ are from (3), symbols show various parameter sets (see the text). (d) Results from DMRG $12\!\times\!12$ cluster calculations for the set (3) with spins in the $yz$ plane.
[…]
@@
[…]
@@
This is the set of parameters that provides a close fit to all available INS data for BaCAO in the polarized phase, as shown in Figs. 1 and 2, and Figs. 3(a) and 3(b) for ${\bf B}\!\parallel\!b$ at B=3B\!=\!3 T with the data from Ref. [40]. In Fig. 3(a), we show the data in multiple $\bf q$}-panels together with the $XXZ$ fit from Ref. [40] (black dashed lines) and our results for the single-magnon branches and the bottom of the two-magnon continuum (solid and light-blue dashed lines, respectively). Fig. 3(b) shows a side-by-side comparison of the intensity maps of the dynamical structure factor ${\cal S}({\bf q},\omega)$ from Ref. [40] and our theoretical results. The comparison of our results with the data for B=0.75B\!=\!0.75 T and ${\bf B}\!\parallel\!b$ from Ref. [40] is shown in EM.
@@
In Fig. 3(c), we present the classical Luttinger-Tisza (LT) ${\sf J}_{z\pm}$${\sf J}_{3}$ phase diagram of the model (1)++(2) for $\Delta_{1}\!=\!0.36$ and ${\sf J}_{2}/{\sf J}_{1}\!=\!0.03$ as in (3), and we set $\Delta_{3}\!=\!{\sf J}_{\pm\pm}\!=\!0$ for simplicity [19, 64]. The sequence of the FM, ZZ, and helical phases shown in Fig. 3(c) is common in many classical honeycomb-lattice models [11, 54, 19].
@@
Recent studies of the quantum S=1/2S\!=\!1/2 edition of the XXZXXZ ${\sf J}_{1}$${\sf J}_{3}$ model using the density-matrix renormalization group (DMRG) method [54, 55] found no trace of the spiral phase in its phase diagram, uncovering instead only narrow ranges of non-classical states, such as dZZ. Here, we extend this study to the ${\sf J}_{1}$${\sf J}_{z\pm}$${\sf J}_{3}$ model, with the remaining parameters ($\Delta_{n}$, $\sf J$$_{\pm\pm}$, and ${\sf J}_{2}$) fixed according to the proposed BaCAO set in (3).
@@
DMRG simulation of the proposed parameter set (3) on a $12\times 12$ cylinder are shown in Fig. 3(d), demonstrating the dZZ state with spins shown in the $yz$ plane to emphasize their out-of-plane tilt. While the non-classical dZZ state for the XXZXXZ ${\sf J}_{1}$${\sf J}_{3}$ models has recently received theoretical endorsements [42, 66, 58, 55], the tilt is unusual and is clearly due to the ${\sf J}_{z\pm}$ coupling. According to our DMRG results, every second spin in the dZZ structure is tilted out of plane by about $35\degree$. Using the g-factors $g_{z}/g_{ab}\!\approx\!0.5$, this yields $7.8\degree$ average tilt of the magnetic moment, compared with $\sim\!6\degree$ suggested by neutron polarimetry [12, 13]. These and other quantitative discrepancies can be attributed either to finite-size effects in DMRG or to small terms in the description of BaCAO that are not accounted for by our model [62].